How many grams of 13% HCl aqueous solution by mass should be taken to prepare 0.5 L 0.646 M HCl solution?
1
Expert's answer
2020-11-20T10:07:05-0500

CM(HCl)=n/V

n (HCl)=CM x V = 0.646 x 0.5 = 0.323 mol

n = m/M

m (HCl) =n x M

M (HCl) = 36.46 g/mol

m (HCl) = 0.323 x 36.46 = 11.78 g


13 g of HCl is contained in 100 g of 13% solution

11.78 g is contained in x g of 13% solution

m(HCl)13%solution=(11.78 x 100)/13 = 90.6 g